create the amortization schuelde for a loan of $4800, paid monthly over 2 years using an apr of 8 percent
Answer #1
Amortization Schedule:
Month end Beginning Loan Interest Expense Monthly Payment Reduction of Loan Ending Loan Amount
a b=a*8%*1/12 c d=c-b e=a-d
1 $ 4,800.00 $       32.00 $     217.09 $     185.09 $ 4,614.91
2     4,614.91           30.77         217.09         186.32     4,428.58
3     4,428.58           29.52         217.09         187.57     4,241.02
4     4,241.02           28.27         217.09         188.82     4,052.20
5     4,052.20           27.01         217.09         190.08     3,862.12
6     3,862.12           25.75         217.09         191.34     3,670.78
7     3,670.78           24.47         217.09         192.62     3,478.16
8     3,478.16           23.19         217.09         193.90     3,284.26
9     3,284.26           21.90         217.09         195.20     3,089.06
10     3,089.06           20.59         217.09         196.50     2,892.56
11     2,892.56           19.28         217.09         197.81     2,694.76
12     2,694.76           17.97         217.09         199.13     2,495.63
13     2,495.63           16.64         217.09         200.45     2,295.18
14     2,295.18           15.30         217.09         201.79     2,093.39
15     2,093.39           13.96         217.09         203.14     1,890.25
16     1,890.25           12.60         217.09         204.49     1,685.76
17     1,685.76           11.24         217.09         205.85     1,479.91
18     1,479.91              9.87         217.09         207.22     1,272.69
19     1,272.69              8.48         217.09         208.61     1,064.08
20     1,064.08              7.09         217.09         210.00         854.08
21         854.08              5.69         217.09         211.40         642.68
22         642.68              4.28         217.09         212.81         429.88
23         429.88              2.87         217.09         214.23         215.65
24         215.65              1.44         217.09         215.65            -0.00
Total $     410.18 $ 5,210.18 $ 5,620.37
Working:
Present value of annuity of 1 = (1-(1+i)^-n)/i Where,
= (1-(1+0.006667)^-24)/0.006667 i = 8%/12 = 0.006667
= 22.110544 n = 2*12 = 24
Monthly Payment = Loan amount / Present value of annuity of 1
= $ 4,800.00 / 22.110544
= $     217.09